# rtboxstyle linecolor=black, backgroundcolor=lightlight-gray, leftmargin=0cm, rightmargin=0cm, skipabove=4pt, skipbelow=4pt, linewidth=0.4pt, innertopmargin=3pt, innerbottommargin=3pt, innerrightmargin=6pt, innerleftmargin=6pt, roundcorner=2

### v1

@@ Appendix C Hyperparameters Setting
Hidden size is of LSTM for heterogeneous graph generation. Hidden size is a key parameter for LSTM, as it determines the capacity of the model to capture temporal and structural dependencies in sequential node and edge generation. We conducted experiments on the Nodlink dataset and evaluated structural fidelity using five MMD-based metrics in the validation set. The results in Figure 4 show that a hidden size of 64 yields significantly lower fidelity compared to the other settings. For hidden sizes of 128, 256, and 512, the MMD values remain at a comparable level, with 256 performing slightly better than 128 and 512. Considering both performance and computational cost, we select 256 as the hidden size for ProvSyn.
Inference temperature is a key parameter in LLM-based inference, controlling the randomness of generated text. Higher temperatures yield more diverse outputs, while lower temperatures produce more deterministic results. We conducted experiments on the Nodlink dataset using temperature values of 0.1, 0.5, 1.0, and 1.5, with all other inference parameters held constant. Evaluation focused on the quality of text generated for process and file node attributes, as well as semantic correctness. As shown in Figure 5, the best performance is achieved when the temperature is set to 1.5. In particular, at lower temperatures, the generated node names tended to be repetitive and lacked diversity. Based on these findings, we selected 1.5 as the final inference temperature for ProvSyn, consistent with the recommendation of the FastLanguageModel library.
Appendix D Ablation Study
Alternative models to GraphGen. In the heterogeneous graph generation module, we adopt the design of GraphGen, which encodes graphs as DFS code sequences and employs an LSTM for training and inference. To validate the effectiveness of this design, we compare it with alternative models, including the widely used GraphRNN [72]. In our experiments, we evaluate the graphs generated by GraphRNN and GraphGen using five MMD metrics. As shown in Figure 6(a), the graphs generated by GraphGen consistently achieve significantly lower MMD scores across all five metrics, indicating substantially better fidelity compared to those generated by GraphRNN. These results demonstrate the effectiveness of using GraphGen for heterogeneous graph generation in ProvSyn.
Masking strategies. For text attribute generation, we employ two masking strategies to construct the training set, corresponding to two different input sequences during inference. To evaluate the effectiveness of the masking design, we perform ablation experiments by disabling one of the strategies and using only the other. Model performance is then evaluated using text quality metrics and semantic correctness. The experiments are conducted on the Trace dataset, with the number of training samples fixed at 20,000 for both the single-strategy and dual-strategy settings. In the results shown in Figure 6(b), we observe that using only the Full Masking strategy yields better performance than using only the Part Masking strategy. This may be because the Part Masking strategy provides partial node names within the sequence, offering more cues for the LLM to leverage. As a result, a model trained solely with Part Masking performs poorly when tested on Full Masking sequences, which contain fewer clues and less information. In contrast, using both strategies together achieves better performance than either alone, as it exposes the model to a broader range of sequence types that may appear during inference. These results demonstrate the effectiveness of our proposed approach that combines both masking strategies.
LLM parameter size. In our experiments on text attribute generation, we employ the Llama-3.2-3B-Instruct model as the base model for training. We also compare its performance against models of different sizes, including Llama-3.2-1B-Instruct and Llama-3.1-8B-Instruct. All models are trained on the same training datasets and evaluated using both text quality metrics and the semantic correctness metric. The results, presented in Figure 6(c), demonstrate that the 3B model achieves the highest scores on the BLEU and GLEU metrics. In contrast, the 8B model shows marginally better performance on the ROUGE and semantic correctness metrics. Given that the advantage of the 8B model is not substantial and considering computational resource constraints, we select the 3B model as the base for synthesizing text attributes. In practice, ProvSyn introduces a general training strategy to leverage LLMs for synthesizing text attributes, where the choice of the base model involves balancing the complexity of the dataset, the application requirements and the limitations of computational resources.

### v2

@@ Appendix E Ablation Study
[Task Description] You are an audit log generator. Carefully analyze the audit log records provided in [Examples] to understand their structure, semantics, and temporal patterns. Based on this understanding, generate 50 unique audit log records that are logically consistent, structurally valid, and realistic. The generated logs should reflect plausible system activities, preserve coherent relationships between entities and events, and follow realistic ordering observed in the examples.
[Examples]
"event_type": "EVENT_OPEN",
"source_node_type": "file",
"source_node_name": "kernel_mem.txt",
"target_node_type": "subject",
"target_node_name": "vmstat"
"event_type": "EVENT_READ",
"source_node_type": "file",
"source_node_name": "None",
"target_node_type": "subject",
"target_node_name": "pickup"
"event_type": "EVENT_WRITE",
"source_node_type": "subject",
"source_node_name": "head",
"target_node_type": "file",
"target_node_name": "top_procs.txt"
"event_type": "EVENT_CONNECT",
"source_node_type": "subject",
"source_node_name": "master",
"target_node_type": "file",
"target_node_name": "None"
"event_type": "EVENT_EXECUTE",
"source_node_type": "file",
"source_node_name": "/usr/bin/top",
"target_node_type": "subject",
"target_node_name": "top"
[Output] Return the logs in JSON format. Only output the JSON.
Figure 7: Prompt for LLM Provenance Data Generation on Cadets-E3